OpenStudy (chudspeth):

It would take Marco 6 hours to paint a fence by himself. It would take Jaylyn 5 hours to paint the same fence by herself. Marco is creating this table to help him determine how long it would take them to paint the fence together. Which heading should be written above the first column of numbers?

OpenStudy (eliesaab):

Let t be the number of hours needed to paint the fence together then in one hour they paint \[\frac 1 t \] fo the fence

OpenStudy (eliesaab):

This should be equal to \[ \frac 1 6 + \frac 1 5 \]

OpenStudy (eliesaab):

Hence \[ \frac 1 t = \frac 1 5 + \frac 1 6 \] Solve for \(t\) and you are done

OpenStudy (eliesaab):

You find \[ t= \frac {30}{11} \]